# Topologically protected flatness in chiral moiré heterostructures

### v1

@@ IV.1 Hierarchy between disorder channels
As represented in Fig. 5, we split the various types of disorder susceptible to broaden the flat bands of chiral TBG into three categories: the off-resonant terms scattering particles from the flat-band to degrees of freedom above or below the spectral gap, the chiral terms that anticommute with the abelian chiral symmetries of the two flat bands (they can only be defined for the first magic angle, see Sec. III), and resonant non-chiral scattering processes directly coupling the two sublattice polarized flat-bands. In this paragraph, we identify what physical mechanism, if any, protects the flat bands against those three types of disorders.
Figure 5: Schematic representation of the different disorder channels considered, e.g. in Tab. 1, using the band structure of chiral TBG at the first magic angle along a high-symmetry line of the moiré Brillouin zone. The two zero-energy sublattice-polarized flat bands have been shifted in energy for visibility.
Off-resonant channels protected by the spectral gap:
@@
Disorder channels of the first kind connect the flat bands to degrees of freedom away from zero energy. Included in this category is, for instance, inter-layer tunneling that preserves the sublattice. Suppose the disorder strength is small compared to the spectral gap separating the flat bands from the rest of the spectrum. In that case, such disorder only contributes to the bandwidth of the flat band to second order, thereby granting an effective protection of the flat band against off-resonant scattering terms. Because the spectral gap separating the flat bands from the rest of the spectrum exponentially decays with $\alpha$ for large $\alpha$ [67], we expect that higher magic angles will be more susceptible to off-resonant disorder than the abelian flat band of the first magic angle. This is especially significant because the decay of the spectral gap is found to be extremely steep: the gap above the flat bands of the three first magic angles $\alpha=0.586,2.221,3.751$ is respectively equal to $0.29$, $0.059$ and $0.0027$ in units of the typical moiré kinetic energy $|v\kappa|$ (that itself decreases linearly with $\theta$). The asymptotic scaling of the spectral gap for large $\alpha$ was recently used to prove the impossibility of observing very small magic angles in chiral TBG [51].
Topologically protected channels:
Another difference between the first and higher magic angles is the presence of an index theorem, derived in Sec. III, which has direct consequences on the robustness of the flat bands against another type of disorder (which can be both off-resonant and resonant). Indeed, the topological nature of the AS theorem ensures perfect immunity of the flat bands of the first magic angle against disorder channels that simultaneously anticommute with the chiral symmetries of the two abelian Dirac operators $\mathcal{D}^{\prime}_{\pm}$ (Sec. II). Because of the $U$ transformation, these two chiral symmetries are not exactly equal to the chiral operator $\sigma_{3}$ of the original Hamiltonian Eq. 3, but to its projection onto the two abelian sectors. Therefore, the perturbations that leave the two flat bands intact must simultaneously anticommute with $\sigma_{3}$ and commute with the constant of motion $\vec{X}\cdot\vec{\mu}$. The only channels of disorder complying with these constraints are $\sigma_{1/2}\mu_{0}$ and $\sigma_{1/2}(\vec{X}\cdot\vec{\mu})$. The former represents a homo-strain that identically couples to both layers. The latter is a fine-tuned combination of interlayer tunneling and heterostrain. In practice, any realistic disorder will contain some non-zero weight onto these two topologically protected channels, further lowering the overall strength of the bandwidth-generating part of the disorder for the first magic angle.
Localization protected channels:
We now turn to resonant scattering channels between the two sublattice-polarized flat bands and first consider those carried by the off-diagonal layer Pauli matrices $\mu_{1,2}$, e.g. higher harmonics of the $AB$ interlayer tunneling. Any inter-layer term of this form preserving $C_{3}$ symmetry must vanish at the $AA$ stacking point. If the disorder only preserves $C_{3}$ on average, and if it does not fluctuate too wildly over the moiré scale (i.e. if the disorder auto-correlation distance is not too small compared to the moiré lattice constant), it will remain strongly suppressed in the $AA$ region. Now, the wavefunctions of the first magic angle flat bands in chiral TBG are known to be strongly peaked around the $AA$ point [68, 69, 70]. Due to additional nodes near the $AA$ region, higher magic angles have their density localized closer to the regions where the inter-layer resonant disorder is strong [66], which can be seen in Fig. 4 (note the logarithmic scale). A direct consequence of this fact is a lower sensitivity of the first magic angle to inter-layer resonant disorder, i.e. an effective protection from the localization of the wavefunction.
Unprotected channels:
At last, we consider resonant intra-layer disorder, corresponding, for instance, to hetero-strain. These disorder channels can have large amplitude near the $AA$ region if they preserve $C_{3}$ on average. The previous argument is therefore reversed, and we now expect higher magic angles having nodes near the $AA$ region to enjoy a lower sensitivity to this disorder. On the other hand, the first magic angle is expected to be very sensitive to hetero-strain. This has indeed been observed numerically, with the landscape of quantum phases stabilized by interactions drastically changing upon including hetero-strain of very weak amplitude [71, 72, 73, 74].
@@
To summarize (see Tab. 1), the first magic angle possesses an intrinsic robustness to a broad class of disorder channels thanks to its topological origin, large gap, and the localization of its wavefunctions. It is, however, sensitive to hetero-strain, which can lead to substantial changes in the quantitative physics of TBG [75]. On the other hand, higher magic angles are expected to strongly broaden in the presence of any disorder except intra-layer resonant channels.
@@ IV.2 Numerical simulations
To solidify these observations, we numerically simulate the effect of disorder in the chiral model of TBG. The role of the AS index in the robustness against disorder can be further highlighted by focusing on disorder in the $AB$ inter-layer tunneling channel, for which all magic angles possess some protection due to the localization of their wavefunctions and where the first magic angle enjoys an additional protection due to its abelian nature (third line of Tab. 1).
To be more concrete, we introduce a $C_{3}$ symmetric disorder by adding higher-order harmonics to the inter-layer tunneling potential. This amounts to the substitution $\mathcal{A}_{z}\to\mathcal{A}_{z}+\delta\mathcal{A}_{z}$ in Eq. 3 with
@@
where $\kappa_{m,0}$ is the $m$-th smallest vector in the first trine of the plane that connects the two corners $\kappa$ and $\kappa^{\prime}$ of the moiré Brillouin zone (up to reciprocal lattice vectors), $\kappa_{m,n}$ denotes its rotation by $32n\pi/$ 3, and $\mu_{m}$ stands for the amplitude of the $m$-th harmonic. The $\phi_{m,n}$ phases are included such that $\delta\mathcal{A}_{z}$ does not break $C_{3}$ and the $C_{2y}$ symmetry of Eq. 3. The precise definition of all these terms is given in App. B. We only consider the first $m_{c}=10$ harmonics and discretize $\mathcal{D}$ on a momentum-space lattice containing all plane waves with wavevectors of norm smaller than a certain radius $R_{k}$.
For the considered $C_{3}$-preserving inter-layer tunneling disorder, the chiral model of TBG still possesses exact magic angles where the bands at zero energy become perfectly degenerate [76, 56]. The precise value of these magic $\alpha$ can be efficiently determined using the method of Refs. [56, 67], where all magic angles are determined as eigenvalues of a carefully designed operator $\Lambda$, whose explicit expression is given in App. B for comprehensiveness. We first use this method to characterize the sensitivity of the magic angles to the different harmonics in Eq. 9. For this purpose, we choose one specific harmonic $m$ and compute the change $\delta\alpha$ in the first, second, and third magic angles when $\mu_{m}=\delta\mu$ is set to a small but finite value (keeping all other $\mu_{n\neq m}=0$). For small disorder strength, this corresponds to taking the derivatives $\partial\alpha/\partial\mu$ of $\Lambda$’s eigenvalues with respect to $\mu_{m}$. We numerically evaluate this derivative using a small enough $\delta\mu=0.02$, and plot our results for $|\partial\alpha/\partial\mu|$ in Fig. 6a. We observe that the $\alpha$ deviations in the first magic angle exponentially decay with $m$ and are negligible for $m>5$, indicating that the first magic angle is robust against microscopic details on scales smaller than the moiré length. In contrast, the deviations for the higher magic angle can reach up to ten times the amplitude of the added perturbation and do not decay with $m$. This suggests that the higher magic angles are easily affected by microscopic disorder at all scales down to graphene’s lattice constant. Perturbations on those scales, such as lattice relaxations, are inevitable and necessarily spoil the prediction for higher magic angles [41, 56]. Perturbations on scales larger than the moiré length are also studied in App. C.
@@
To understand the physical implications of the difference in $\delta\alpha$ for first and higher magic angles observed in Fig. 6a, we consider realistic disorder realization containing all harmonics $m\leq m_{c}$ with amplitudes $\mu_{m}$ drawn uniformly in $[-1,1]$ with the overall normalization $\sqrt{\sum_{m}\mu_{m}^{2}}=\delta\mu=0.1$. We plot the distribution of the first three magic angles for 500 realizations of such disorder in Fig. 6b, which shows a similar trend as Fig. 6a: the distribution of magic $\alpha$ is sharply peaked around the original value 0.586 for the first magic angle. At the same time, it is extremely broad and featureless for higher magic angles. To interpret this result, picture a sample in which different patches of the sample (different unit cells) are characterized by slightly different atomic arrangements and lattice relaxation profiles that locally change the values of the $\mu_{m}$. We can assign an effective magic angle $\alpha$ to each of these regions. When TBG is close to the first unperturbed magic angle, the distribution of these magic $\alpha$ is peaked, and the magic condition is almost simultaneously satisfied for all patches (see Fig. 6b). This is not true for higher magic angles, such that the density of states at zero energy is smeared in many patches, yielding a strong broadening of the original flat bands.
We can estimate the flat band broadening by computing $E_{\gamma}$ the average energy of the topmost band at the center of the moiré Brillouin zone ($\gamma$ point) for 500 realizations of the same realistic disorder pattern for a fixed twist angle, i.e. as a function $\alpha$. The results presented in Fig. 6c pictorially capture our previous statement: the higher magic angles are completely washed out due to their high sensitivity to disorder. The average $\gamma$-point energy only displays an apparent non-monotonic behavior with an almost vanishing bandwidth at the first magic angle, which we have seen to be protected against disorder (Tab. 1).
Our numerical calculations clearly show that only the flat bands at the first magic angle retain a large density of states in the presence of disorder. As explained in Tab. 1, this is due to the combination of the topological protection from the AS theorem, large spectral gap, and strong localization of the wavefunctions in the flat band. Higher magic angles are sensitive to lattice scale details and should be interpreted as accidental artifacts of the continuum model rather than experimentally observable features.
@@ IV.3 Away from the chiral limit
We finally check whether our results, which find their roots in the chiral limit of TBG, also hold in the more realistic situation where non-zero $AA$ tunnelings are included in the Bistritzer-MacDonald model [40]. When the $AA$ tunneling amplitude $w_{0}$ is non-zero, we cannot identify the magic angle with the point where the $\gamma$-point energy vanishes, as we did in Fig. 6c, since the model never features perfectly flat bands. Instead, we locate the magic angle as the points for which the Fermi velocity at the corner of the moiré Brillouin zone vanishes [76, 40]. We have numerically computed this Fermi velocity $v_{m}(\alpha)$ averaged over 500 realizations of the generic disorder pattern described above as a function of $\alpha$, both in the chiral limit $w_{0}=0$ and for the realistic value $w_{0}=0.8w$ [40]. Our results are presented in Fig. 7, where, for better comparison, we have chosen $\delta\mu=0.1$ for the chiral limit and $\delta\mu=0.2$ in the realistic regime, leading to the similar average velocity at the first magic angle. In both the chiral and realistic (nonchiral) limits, disorder barely affects the Fermi velocity (blue lines) from its clean values (dotted lines) as long as we are in the abelian regime ($\alpha<0.86$). In the non-abelian regime, the average velocity deviations become significant. Consequently, only the first magic angle still strongly suppresses the Fermi velocity $v_{m}(\alpha)$. In contrast, all higher magic angles are washed out by disorder – we interpret this as a direct consequence of the robustness granted by the AS topological index described in the second line of Tab. 1.

### v2

@@ IV.1 Hierarchy between disorder channels
We split the various types of disorder susceptible to broadening the flat bands of chiral TBG into three categories: the chiral channels anticommuting with the chiral operators of the two decoupled abelian components of the Hamiltonian, the resonant scattering processes directly coupling the two sublattice polarized flat-bands, and the off-resonant terms scattering particles from the flat-band to degrees of freedom above or below the spectral gap. We here provide a physical discussion of the effects of these different disorder channels on the flat bands.
IV.1.1 Chiral disorder
The Atiyah-Singer index ensures the stability of the flat bands against any perturbations anticommuting with the chiral symmetry $\sigma_{3}$ in each of the decoupled abelian components of the Hamiltonian separately. In other words, after the basis transformation of Eq. 4, the system can sustain any kind of disorder in the $\sigma_{1/2}\mu_{0}$ and $\sigma_{1/2}\mu_{3}$ channels. The first of these operators is independent of the layer-basis and takes the same form in the original degree of freedom of the problem. In the physical basis of Eq. 3, disorder in the $\sigma_{1/2}\mu_{0}$ channel describes homo-strain, against which all the magic angle are protected by virtue of the AS index theorem – as summarized in the first line of Tab. 1.
In contrast, transforming $\sigma_{1/2}\mu_{3}$ back to the original basis gives
$P(\vec{X})=\begin{bmatrix}0&\vec{X}\cdot\vec{\mu}\\ \vec{X}^{*}\cdot\vec{\mu}&0\end{bmatrix},$ (15)
a fine-tuned disorder channel that explicitly depends on $\vec{X}$, and hence on the details of the abelianization at each magic angle. While this disorder channel contains fine-tuned combinations of layer Pauli operators, it can still efficiently protect the band flatness provided realistic forms of disorder largely overlap with it. To determine whether typical disorder realization can, in principle, overlap with $P(\vec{X})$, it is instructive to look more closely at the structure of the constant of motion $\vec{X}$ at the various magic angles. In Fig. 3, we display $||X(q)||=[\sum_{i}|X_{i}(q)|^{2}]^{1/2}$ the norm of the Fourier transform of $\vec{X}$. We observe that it is sharply peaked at $q=0$ for the first magic angle, highlighting that $P(\vec{X})$ acts locally, whereas $||X(q)||$ exhibits maxima at finite momenta for higher magic angles. As a result, random disorder realization carrying no zero average momentum will only carry high overlap with $P(\vec{X})$ at the first magic angle. This suppression of the disorder strength due to finite overlap with $P(\vec{X})$ concerns all channels contained in it and includes both hetero-strain and variations of the $AB$-tunneling amplitude.
Figure 3: Norm $||X(q)||=[\sum_{i}|X_{i}(q)|^{2}]^{1/2}$ in momentum space for the first (blue), second (orange), and third (orange) magic angles. Only at the first magic angle is $P(\vec{X})$ dominated by its $q=0$ component and therefore acts locally. For visibility, we only included momenta with $q_{x},q_{y}\geq 0$.
IV.1.2 Non-chiral resonant disorder
Let us now consider disorder channels that do not necessarily anti-commute with the chiral symmetries in the decoupled abelian basis, but nevertheless yield direct coupling between the two sublattice flat bands. These disorder channels are carried by the $\sigma_{1/2}$ orbital Pauli matrices and hence anticommute with the original chiral operator $\sigma_{3}$ of the original chiral Hamiltonian of TBG. For the sake of concreteness, we focus on the $\sigma_{1/2}\mu_{1/2}$ channels, which physically correspond to perturbations to the $AB$ tunneling amplitude that modify the effective gauge potential $\mathcal{A}_{z}$ by introducing higher moiré harmonics and can be due, for instance, to out-of-plane lattice relaxation.
These resonant disorder channels induce a first-order broadening of the band proportional to a matrix element connecting the two sublattice polarized flat bands by some high moiré harmonics. This matrix element is small (resp. large) if the dominant disorder harmonics is large (resp. small) compared to the extent of the flat band wavefunctions in momentum space, such that the disorder matrix element involves momentum points with weak (resp. large) amplitude of wavefunctions (App. C). Indeed, the envelope of the flat band wavefunctions $|\psi_{A,k}\rangle$ in momentum space is determined by the transformation $U$, which at the magic angle is proportional to $\vec{X}\cdot\vec{\mu}$ (see Sec. III.4). The functional form of the $\vec{X}$ vector therefore strongly influences the effects that disorder has on the flat bands. As a result, the flatness of the band is protected against most forms of tunneling disorder provided $\vec{X}$ is strongly localized in momentum space, and strongly susceptible to broadening when $\vec{X}$ has support over a large number of harmonics in momentum space. In Fig. 3, we show that $\vec{X}$ is strongly localized in momentum space only at the first magic angle, which again highlights the importance of the locality of the constant of motion $\vec{X}$ in the mitigation of disorder and protection of the band flatness. This is summarized in the second line of Tab. 1.
More formally, the magic angles can be obtained from the eigenvalues of the operator $\Lambda=\hat{D}^{-1}\hat{T}_{+}\hat{D}^{-1}\hat{T}_{-}$ (see App. C[55], where $T_{\pm}=\mathcal{A}_{z}(\pm r)$ and $\hat{D}^{-1}$ is the Dirac propagator. Disorder modifies the tunneling terms as $T\rightarrow T+\delta T$. We can now use perturbation theory in $\delta T$ to understand the robustness of the magic angle. The perturbation $\delta T$ transfers momenta from $k$ to $k+q$, and the propagator penalizes this transfer by a factor $1/(k+q)$. As a result, for a tunneling disorder with correlation length $\xi$, wavefunctions that are confined in momentum-space, with size $k_{\xi}\sim 2\pi/\xi$, are less prone to broadening. As shown in the App. D but also explicit in Fig. 2d-f, the wavefunction of the first magic angle is the most confined in momentum space and is, hence, the most robust. In our numerical simulation presented below, we find that the position of the first magic angle remains remarkably robust against these perturbations, while the position of higher magic angles is extremely sensitive to their effects.
IV.1.3 Non-chiral non-resonant disorder
@@
Disorder channels of this kind connect the flat bands to degrees of freedom away from zero energy. Included in this category is, for instance, inter-layer tunneling that preserves the sublattice. If the disorder strength is small compared to the spectral gap between the flat bands and the rest of the spectrum, the disorder only affects the flat band’s bandwidth in second order. This provides an effective protection of the flat band against off-resonant scattering.
In Ref. [69], it was shown that the spectral gap between the flat bands and the rest of the spectrum decays exponentially as a function of $\alpha$ as the latter increases. Using the decoupling introduced in Sec. III, we interpret this as a result of the exponential squeezing of the spectrum of Landau-like operators as a function of the amplitude of the magnetic field’s fluctuations [56, 57, 58, 70]. Fig. 2d-f illustrate that these fluctuations grow rapidly with $\alpha$: they increase more than tenfold with each successive magic angle, and are only smaller than the average magnetic field at the first magic angle. This rapid increase leads to a sharp collapse of the spectral gap above the flat bands: at the first three magic angles $\alpha=0.586,2.221,3.751$ the spectral gap respectively read $0.29$, $0.059$ and $0.0027$ in units of the typical moiré kinetic energy $|v\kappa|$ (that itself decreases linearly with $\theta$). The increasing relevance of the non-uniform part of the magnetic field, leading to an exponential reduction of the gap as a function $\alpha$ for large $\alpha$, was recently used to prove the impossibility of observing higher magic angles in TBG [70]. This is summarized in the last line of Tab. 1.
@@
To summarize (see Tab. 1), the first magic angle possesses intrinsic robustness to a broad class of disorder channels thanks to its topological origin, large gap, and the localization of its wavefunctions that originates from its well-behaved constant of motion $\vec{X}$. On the other hand, higher magic angles are expected to strongly broaden in the presence of any disorder except homo-strain. We now probe this qualitative difference numerically.
@@ IV.2 Numerical simulations
To bolster these arguments, we numerically simulate the effect of disorder in the chiral model of TBG. Concretely, we introduce a $C_{3}$ symmetric disorder by adding higher-order harmonics to the inter-layer tunneling potential. This amounts to the substitution $\mathcal{A}_{z}\to\mathcal{A}_{z}+\delta\mathcal{A}_{z}$ in Eq. 3 with
@@
where $\kappa_{m,0}$ is the $m$-th smallest vector in the first trine of the plane that connects the two corners $\kappa$ and $\kappa^{\prime}$ of the moiré Brillouin zone (up to reciprocal lattice vectors), $\kappa_{m,n}$ denotes its rotation by $32n\pi/$ 3, and $\mu_{m}$ stands for the amplitude of the $m$-th harmonic. The $\phi_{m,n}$ phases are included such that $\delta\mathcal{A}_{z}$ does not break $C_{3}$ and the $C_{2y}$ symmetry of Eq. 3. The precise definition of all these terms is given in App. D. We only consider the first $m_{c}=10$ harmonics and discretize $\mathcal{D}$ on a momentum-space lattice containing all plane waves with wavevectors of norm smaller than a certain radius $R_{k}$.
For the considered $C_{3}$-preserving inter-layer tunneling disorder, the chiral model of TBG still possesses exact magic angles where the bands at zero energy become perfectly degenerate [71, 55]. The precise value of these magic $\alpha$ can be efficiently determined using the method of Refs. [55, 69], where all magic angles are determined as eigenvalues of the operator $\Lambda$ (see App. D for details). We first use this method to characterize the sensitivity of the magic angles to the different harmonics in Eq. 16. For this purpose, we choose one specific harmonic $m$ and compute the change $\delta\alpha$ in the first, second, and third magic angles when $\mu_{m}=\delta\mu$ is set to a small but finite value (keeping all other $\mu_{n\neq m}=0$). For small disorder strength, this corresponds to taking the derivatives $\partial\alpha/\partial\mu$ of the eigenvalues of $\Lambda$ with respect to $\mu_{m}$. We numerically evaluate this derivative using a small enough $\delta\mu=0.02$, and plot our results for $|\partial\alpha/\partial\mu|$ in Fig. 4a. We observe that the $\alpha$ deviations in the first magic angle exponentially decay with $m$ and are negligible for $m>5$, indicating that the first magic angle is robust against microscopic details on scales smaller than the moiré length. In contrast, the deviations for the higher magic angle can reach up to ten times the amplitude of the added perturbation and do not decay with $m$. This suggests that the higher magic angles are easily affected by microscopic disorder at all scales down to graphene’s lattice constant. Perturbations on those scales, such as lattice relaxations, are inevitable and necessarily spoil the prediction for higher magic angles [43, 55]. Perturbations on scales larger than the moiré length are also studied in App. D.
@@
To understand the physical implications of the difference in $\delta\alpha$ for first and higher magic angles observed in Fig. 4a, we consider realistic disorder realization containing all harmonics $m\leq m_{c}$ with amplitudes $\mu_{m}$ drawn uniformly in $[-1,1]$ with the overall normalization $\sqrt{\sum_{m}\mu_{m}^{2}}=\delta\mu=0.1$. We plot the distribution of the first three magic angles for 500 realizations of this disorder in Fig. 4b, which shows a similar trend as Fig. 4a: the distribution of magic $\alpha$ is sharply peaked around the original value 0.586 for the first magic angle. At the same time, it is extremely broad and featureless for higher magic angles. To interpret this result, picture a sample in which different patches of the sample (different unit cells) are characterized by slightly different atomic arrangements and lattice relaxation profiles that locally change the values of the $\mu_{m}$. We can assign an effective magic angle $\alpha$ to each of these regions. When TBG is close to the first unperturbed magic angle, the distribution of these magic $\alpha$ is peaked, and the magic condition is almost simultaneously satisfied for all patches (see Fig. 4b). This is not true for higher magic angles, such that the density of states at zero energy is smeared in many patches, yielding a strong broadening of the original flat bands.
We can estimate the flat band broadening by computing $E_{\gamma}$ the average energy of the topmost band at the center of the moiré Brillouin zone ($\gamma$ point) for 500 realizations of the realistic disorder pattern of Eq. 16 for a fixed twist angle, i.e. as a function $\alpha$. The results presented in Fig. 4c pictorially capture our previous statement: the higher magic angles are completely washed out due to their high sensitivity to disorder. The average $\gamma$-point energy only displays an apparent non-monotonic behavior with an almost vanishing bandwidth at the first magic angle, which we have seen to be protected against disorder (Tab. 1).
Our numerical calculations clearly show that only the flat bands at the first magic angle retain a large density of states in the presence of disorder. As explained in Tab. 1, this is due to the combination of the topological protection from the AS theorem to local chiral symmetric disorder, large spectral gap, and strong localization of the wavefunctions in the flat band. Higher magic angles are sensitive to lattice scale details and should be interpreted as accidental artifacts of the continuum model rather than experimentally observable features.
@@ IV.3 Away from the chiral limit
We finally check whether our results, which find their roots in the chiral limit of TBG, also hold in the more realistic situation where non-zero $AA$ tunnelings are included in the Bistritzer-MacDonald model [42]. When the $AA$ tunneling amplitude $w_{0}$ is non-zero, we cannot identify the magic angle with the point where the $\gamma$-point energy vanishes, as we did in Fig. 4c, since the model never features perfectly flat bands. Instead, we locate the magic angle as the points for which the Fermi velocity at the corner of the moiré Brillouin zone vanishes [71, 42]. We have numerically computed this Fermi velocity $v_{m}(\alpha)$ averaged over 500 realizations of the generic disorder pattern described above as a function of $\alpha$, both in the chiral limit $w_{0}=0$ and for the realistic value $w_{0}=0.8w$ [72]. Our results are presented in Fig. 5, where, for better comparison, we have chosen $\delta\mu=0.1$ for the chiral limit and $\delta\mu=0.2$ in the realistic regime, leading to the similar average velocity at the first magic angle. In both the chiral and realistic (nonchiral) limits, disorder barely affects the Fermi velocity (blue lines) from its clean values (dotted lines) as long as we are in the abelian regime ($\alpha\lesssim 1$). In the non-abelian regime, the average velocity deviations become significant. Consequently, only the first magic angle still strongly suppresses the Fermi velocity $v_{m}(\alpha)$. In contrast, all higher magic angles are washed out by disorder – we interpret this as a direct consequence of the robustness granted by the AS topological index described in the second line of Tab. 1.